40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 27     Entry time: Thu Jun 30 16:18:28 2016
Author: Gabriele 
Type: Electronics 
Category: Daily Progress 
Subject: QPD electronics is working 

Today I gave up trying to fix the first board I populated, and built a second one. The good news is that it's working as expected.

With 27.5 uW incident on each quadrant, I measure about 4.5 V, which is in line with the transimpedance of 200k, a responsivity of about 0.4 A/W and ad additional gain of two coming from the differential driver.

I also measured the noise with a SR785 (it wasn't connected to a GPIB interface and I couldn't find any, so all I have are the following numbers and the attached screenshots). 

Frequency [kHz] Noise [uV/rHz] Equivalent beam motion [m/rHz]
1 1.3 5.7e-14
5 2.2 9.6e-14
10 3.3 14e-14
30 2.4 10e-14
60 1.2 5.2e-14

At low frequency we are dominated by 60 Hz harmonics (probably coming from the laser). At high frequency there are some large peaks of unknown origin. I can't acquire digitally the signals to compute the difference, so I don't know if the noise we see is, for example, laser intensity noise. As soon as the cymac is up and running, I'll run some more tests.

 

ELOG V3.1.3-